OpenStudy (anonymous):

help!!!! i think its A but im not sure!!! Christina keeps her DVD collection in a plastic tub. The table shows the number of each type of DVD in Christina's collection. Christina's DVD Collection Type of DVD Number of DVDs Sci-Fi 2 Action 6 Drama 8 Comedy 5 Romance 9 If two DVDs are chosen randomly one at a time, where the first DVD is replaced before the second DVD is chosen, what is the theoretical probability that both DVDs chosen are from the action category? A.1/25 B.1/29 C.1/30 D.6/145

Directrix (directrix):

A.1/25 Of the 30 DVDs, six are action. On the first draw, the probability of selecting an action DVD is 6/30 = 1/5. Because the DVD drawn on the first selection is tossed back into the bin, the probability of selecting an action DVD on the second draw is the same. P( 2 Action DVDs) = 1/5 * 1/5 = 1/25

OpenStudy (anonymous):

thanks so much!!!!

Directrix (directrix):

Glad to help.